OpenStudy (anonymous):

[4.02]What is the y−coordinate of the y−intercept of the line that passes through the points (−4, −4) and (4, 8)?

OpenStudy (anonymous):

First you need to find the slope of the line and then use point slope form

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You got it now? Or need me to work it out

OpenStudy (anonymous):

could you work out

OpenStudy (anonymous):

Yep. What did you get for the slope?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

You should've got 12/8 which simplifies to 3/2... Slope= (y2-y1)/(x2-x1) You got the top and bottom mixed up

OpenStudy (anonymous):

thank you i understand that part and see my error but how do i find the y coordinate

OpenStudy (anonymous):

Ok, use point slope form: y-y1=m(x-x1) pick whichever point you want for (x1, y1) m=slope plug those in

OpenStudy (anonymous):

y-8=3/2(x-4) y-8 = 3/2x - 6 y = 3/2x + 2 2 is the y-coordinatae of the y-intercept